# Real-Time Recurrent Learning using Trace Units in Reinforcement Learning

### v1

@@ Abstract
Recurrent Neural Networks (RNNs) are used to learn representations in partially observable environments. For agents that learn online and continually interact with the environment, it is desirable to train RNNs with real-time recurrent learning (RTRL); unfortunately, RTRL is prohibitively expensive for standard RNNs. A promising direction is to use linear recurrent architectures (LRUs), where dense recurrent weights are replaced with a complex-valued diagonal, making RTRL efficient. In this work, we build on these insights to provide a lightweight but effective approach for training RNNs in online RL. We introduce Recurrent Trace Units (RTUs), a small modification on LRUs that we nonetheless find to have significant performance benefits over LRUs when trained with RTRL. We find RTUs significantly outperform other recurrent architectures across several partially observable environments while using significantly less computation.
@@ 1 Introduction
Agents, animals, and people perceive their surrounding environment through imperfect sensory observations. When the state of the environment is partially observable, agents must construct and maintain their own state from the stream of observations. The constructed agent state summarizes past environment-agent interactions in a form that is useful to predict and control future interactions [40]. Recurrent Neural Networks (RNNs) provide a flexible architecture for constructing agent state [18, 20, 12, 5, 9].
@@
While standard RNN architectures have been mainly supplanted by Transformers [43], in online reinforcement learning settings where the agent learns while interacting with the environment, RNNs remain a promising direction to pursue [16, 10]. There are two main issues that limit the use of self-attention mechanisms from Transformers in online learning. First, calculating the similarity between each pair of points results in a computational complexity that is a function of $k^{2}$, where $k$ is the sequence length. Moreover, calculating the similarity between all pairs ignores the temporal order of the data points, which limits the usefulness of self-attention when the data is temporally correlated [46]. Second, we need access to the whole sequence of observations before taking an action or updating the learnable parameters, which is impractical in continual learning. While recent works have reduced the complexity of transformers from quadratic in the sequence length to linear [14, 15, 36], the full sequence length is still needed to train such architectures. Gated Transformer-XL is an attempt to overcome this issue by keeping a moving window of previous observations [34]. A window of past observations does not scale well to long sequences—the computation is quadratic in the sequence length—and a window is one particular fixed function of history. The simpler recursive form in RNNs, on the other hand, can learn a great variety of functions of history from the data and is well suited for updating the state online for a sequential data stream and have been shown to outperform transformers in such settings [22].
@@
A key open question is how to efficiently train RNNs in online RL. We can divide the literature into methods that approximate Real-Time Recurrent Learning (RTRL) and those that restrict the recurrent architecture. RTRL [45] exploits the recursive nature of the gradient for RNNs, carrying forward the needed gradient information instead of unrolling the recurrent dynamics back in time like Truncated Backpropagation Through Time (T-BPTT) [44]. RTRL avoids storing past data but is so computationally expensive as to be intractable for even moderately sized networks. Several methods approximate the RTRL gradient update, including NoBackTrack [32], Unbiased Online Recurrent Optimization [42, 4], Sparse N-Step Approximation (SnAp) [27]. All of these methods produce a biased gradient estimate. Other works have tried to approximate an unbiased gradient estimate of BPTT specifically for the case of policy gradient updates in RL. However, the approximation resulted in a high variance due to added stochasticity to the policy [1].
@@
In this work, we extend the insights from LRUs into the online RL setting. Our primary contribution are our experiments showing that such a lightweight recurrent architecture can outperform standard approaches like GRUs in RL, with significantly less computation. To obtain this result, we propose a small extension on LRUs, which we call Recurrent Trace Units (RTUs). RTUs incorporate nonlinearity into the recurrence and use a slightly different parameterization than LRUs, but one we find is more amenable to the use of RTRL in online RL than LRUs. We extend Proximal Policy Optimization (PPO) [39] to use RTRL, ablating the decision choices we propose. We provide an in-depth study in an animal-learning prediction benchmark, showing that RTUs scale better than GRUs with increasing compute and number of parameters and that RTUs outperform alternative diagonal recurrent architectures trained with RTRL. We then show across numerous control environments that RTUs have comparable or better performance, compared to GRUs and LRUs.
@@ 2 Background
We formalize the problem setting as a Partially Observable Markov Decision Process (POMDP). At each time step $t=0,1,2,\cdots$, the agent perceives an observation $\mathbf{x}_{t}$, a limited view of the state $\mathbf{s}_{t}\in\mathcal{S}$, and takes an action $A_{t}\in\mathcal{A}(\mathbf{s}_{t})$. Depending on the action taken, the agent finds itself in a new state $\mathbf{s}_{t+1}\in\mathcal{S}$, observes the corresponding observation $\mathbf{x}_{t+1}$ and a reward $R_{t+1}\in\operatorname{\mathbb{R}}$. In the online control setting, the agent’s goal is to maximize the sum of received rewards. It may also make predictions about its environment, such as future observations’ outcomes.
@@
Williams and Zipser (1989) introduced the Real-time Recurrent Learning algorithm (RTRL) as a learning algorithm for continual recurrent learning. Instead of unrolling the recurrent dynamics back in time, RTRL computes the gradient using the most recent observation, and the gradient is calculated and carried from the last step [45]. Assuming the network parameters have not changed, this recursive form gives the exact gradient and does not suffer from the truncation bias inherent to T-BPTT. We provide a more detailed background on the BPTT and RTRL in Appendix B. In reality, the agent updates its parameters frequently, so the gradient information saved from previous time steps is stale, i.e., calculated w.r.t old parameters; yet, under the assumption of small learning rates, RTRL is known to converge [45]. These properties make RTRL ideal for online learning, but unfortunately, there is a catch: it is quartic in the size of $\mathbf{h}_{t}$, which can be prohibitively expensive. For this reason, we pursue a restricted diagonal form in this work, for which RTRL is efficient and linear in $\mathbf{h}_{t}$.
@@ 3.1 Revisiting Complex-valued Diagonal Recurrence
Since we did not impose constraints on the matrix $\mathbf{W}_{h}$, other than being diagonalizable, the eigenvalues of $\mathbf{W}_{h}$ can be complex or real numbers. Previous diagonal RNNs such as eLSTM [16], Columnar networks [17], and IndRNN [19] use only real-valued diagonal matrices. Having only real-valued diagonals assumes that the matrix $\mathbf{W}_{h}$ is symmetric. We provide a small experiment in Appendix A.3 showing that this assumption does not hold even in the simplest setting and that complex eigenvalues do arise. We also investigate whether this result can be extended beyond linear recurrence, and largely obtain a negative theortical result (see Appendix A.1 and  A.2 ).
@@ 3.2 The RTU Parameterization
Each element of $\mathbf{h}_{t}=\mathbf{\Lambda}\mathbf{h}_{t-1}+\mathbf{W}_{x}\ \mathbf{x}_{t}\in\mathbb{R}^{2n}$ has two components $\mathbf{h}_{t}^{c_{1}},\mathbf{h}_{t}^{c_{2}}$, updated recursively
@@
We also adopt two parameterization choices made in LRUs that showed improved performance. The first is learning logarithmic representations of the parameters rather than learning them directly: instead of learning $\mathbf{r}$ and $\boldsymbol{\theta}$, the network learns $\boldsymbol{\nu}^{\log}$ and $\boldsymbol{\theta}^{\log}$, where $\mathbf{r}\doteq\exp(-\boldsymbol{\nu})$, $\boldsymbol{\nu}=\exp(\boldsymbol{\nu}^{\log})$, and $\boldsymbol{\theta}^{\log}\doteq\log(\boldsymbol{\theta})$. This re-parametrization restricts the $\mathbf{r}$ to be $\in(0,1]$, required for stability. We found these modifications to improve stability of RTUs (see Appendix C). The second parameterization choice we adopt from LRUs is to multiply by the input $(\mathbf{W}_{x}\mathbf{x}_{t})_{k}$ with a normalization factor of $\gamma_{k}=(1-r_{k}^{2})^{1/2}$. Putting this all together, the final formulation of RTUs is
$\mathbf{h}^{c_{1}}_{t}$ $=\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}-\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{1}}\mathbf{x}_{t}.$
$\mathbf{h}^{c_{2}}_{t}$ $=\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{2}}\mathbf{x}_{t}.$ (2)
$\mathbf{h}_{t}$ $=[\mathbf{f}(\mathbf{h}_{t}^{c_{1}});\mathbf{f}(\mathbf{h}_{t}^{c_{2}})].$
@@
There are two ways to incorporate non-linearity into RTUs: inside the recurrence or after the recurrence. In the above, in Equation (2), the non-linearity is after the recurrence. These RTUs maintain the equivalence to a dense linear RNN, because the recurrence itself remains linear. We refer to this definition of RTUs as Linear RTUs, because the recurrence is linear, even though we have the ability to represent nonlinear functions by allowing for any nonlinear activation after the recurrence. We also evaluated a different variation of RTUs where the non-linearity is added to the recurrence directly. These Nonlinear RTUs are written as
$\mathbf{h}^{c_{1}}_{t}$ $=\mathbf{f}(\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}-\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{1}}\mathbf{x}_{t})$
$\mathbf{h}^{c_{2}}_{t}$ $=\mathbf{f}(\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{2}}\mathbf{x}_{t})$ (4)
@@ 3.3 The RTRL Update for RTUs
This section shows the RTRL updates for RTUs with more in-depth derivations in Appendix C. To keep notation simpler, we write the updates as if we are directly updating $r$ and $\theta$; the updates for $\boldsymbol{\nu}^{\log}$ and $\boldsymbol{\theta}^{\log}$ are easily obtained then using the chain rule. The full derivation is in Appendix C.2.
Consider the partial derivative with respect to $r_{1}$ for the first RTU with input $\bar{x}_{1}\doteq(\mathbf{W}_{x}^{c_{1}}\mathbf{x}_{t})_{1}$
$h_{t,1}^{c_{1}}=r_{1}\cos(\theta_{1})h_{t-1,1}^{c_{1}}-r_{1}\sin(\theta_{1})h_{t-1,1}^{c_{2}}+\sqrt{(1-r_{1}^{2})}\bar{x}_{1}$
@@
$\frac{\partial\mathcal{L}_{t}}{\partial r_{1}}=\frac{\partial\mathcal{L}_{t}}{\partial h_{t,1}^{c_{1}}}\frac{\partial h_{t,1}^{c_{1}}}{\partial r_{1}}+\frac{\partial\mathcal{L}_{t}}{\partial h_{t,1}^{c_{2}}}\frac{\partial h_{t,1}^{c_{2}}}{\partial r_{1}}$
because $r_{1}$ only impacts the two units in the first RTU, and derivatives w.r.t. the remaining hidden units are zero. Therefore, we just need to keep track of the vector of partial derivatives for these two components, $\mathbf{e}^{r,c_{1}}_{t}\doteq[\frac{\partial h_{t,1}^{c_{1}}}{\partial r_{1}},\ldots,\frac{\partial h_{t,n}^{c_{1}}}{\partial r_{n}}]$ and $\mathbf{e}^{r,c_{2}}_{t}\doteq[\frac{\partial h_{t,1}^{c_{2}}}{\partial r_{1}},\ldots,\frac{\partial h_{t,n}^{c_{2}}}{\partial r_{n}}]$ with recursive formulas

### v2

@@ Abstract
Recurrent Neural Networks (RNNs) are used to learn representations in partially observable environments. For agents that learn online and continually interact with the environment, it is desirable to train RNNs with real-time recurrent learning (RTRL); unfortunately, RTRL is prohibitively expensive for standard RNNs. A promising direction is to use linear recurrent architectures (LRUs), where dense recurrent weights are replaced with a complex-valued diagonal, making RTRL efficient. In this work, we build on these insights to provide a lightweight but effective approach for training RNNs in online RL. We introduce Recurrent Trace Units (RTUs), a small modification on LRUs that we nonetheless find to have significant performance benefits over LRUs when trained with RTRL. We find RTUs significantly outperform other recurrent architectures across several partially observable environments while using significantly less computation. [^1] Code available at https://github.com/esraaelelimy/rtus
@@ 1 Introduction
Agents, animals, and people perceive their surrounding environment through imperfect sensory observations. When the state of the environment is partially observable, agents construct and maintain their own state from the stream of observations. The constructed agent state summarizes past environment-agent interactions in a form that is useful to predict and control future interactions [41]. Recurrent Neural Networks (RNNs) provide a flexible architecture for constructing agent state [19, 21, 13, 6, 10].
@@
While standard RNN architectures have been mainly supplanted by Transformers [44], in online reinforcement learning settings where the agent learns while interacting with the environment, RNNs remain a promising direction to pursue [17, 11]. There are two main issues that limit the use of self-attention mechanisms from Transformers in online learning. First, calculating the similarity between each pair of points results in a computational complexity that is a function of $k^{2}$, where $k$ is the sequence length. Moreover, calculating the similarity between all pairs ignores the temporal order of the data points, which limits the usefulness of self-attention when the data is temporally correlated [47]. Second, we need access to the whole sequence of observations before taking an action or updating the learnable parameters, which is impractical in continual learning. While recent works have reduced the complexity of transformers from quadratic in the sequence length to linear [15, 16, 37], the entire sequence length is still needed to train such architectures. Gated Transformer-XL attempts to overcome this issue by keeping a moving window of previous observations [35]. A window of past observations does not scale well to long sequences—the computation is quadratic in the sequence length—and a window is one particular fixed function of history. The simpler recursive form in RNNs, on the other hand, can learn a great variety of functions of history from the data and is well suited for updating the state online from a sequential data stream and have been shown to outperform transformers in such settings [23].
@@
A key open question is how to efficiently train RNNs in online RL. We can divide the literature into methods that approximate Real-Time Recurrent Learning (RTRL) and those that restrict the recurrent architecture. RTRL [46] exploits the recursive nature of the gradient for RNNs, carrying forward the needed gradient information instead of unrolling the recurrent dynamics back in time like Truncated Backpropagation Through Time (T-BPTT) [45]. RTRL avoids storing past data but is so computationally expensive and is intractable for even moderately sized networks. Several methods approximate the RTRL gradient update, including NoBackTrack [33], Unbiased Online Recurrent Optimization [43, 5], Sparse N-Step Approximation (SnAp) [28]. All of these methods produce a biased gradient estimate. Other works have tried to approximate an unbiased gradient estimate of BPTT specifically for the case of policy gradient updates in RL. However, the approximation resulted in a high variance due to added stochasticity to the policy [1].
@@
In this work, we extend the insights from LRUs into the online RL setting. Our primary contribution are our experiments showing that such a lightweight recurrent architecture can outperform standard approaches like Gated Recurrent Units (GRUs) [4] in RL, with significantly less computation. To obtain this result, we propose a small extension on LRUs, which we call Recurrent Trace Units (RTUs). RTUs incorporate nonlinearity into the recurrence and use a slightly different parameterization than LRUs, but one we find is more amenable to the use of RTRL in online RL than LRUs. We extend Proximal Policy Optimization (PPO) [40] to use RTRL, ablating the decision choices we propose. We provide an in-depth study in an animal-learning prediction benchmark, showing that RTUs scale better than GRUs with increasing compute and number of parameters and that RTUs outperform alternative diagonal recurrent architectures trained with RTRL. We then show across numerous control environments that RTUs have comparable or better performance, compared to GRUs and LRUs.
@@ 2 Background
We formalize the problem setting as a Partially Observable Markov Decision Process (POMDP). At each time step $t=0,1,2,\ldots$, the agent perceives an observation $\mathbf{x}_{t}$, a limited view of the state $\mathbf{s}_{t}\in\mathcal{S}$, and takes an action $A_{t}\in\mathcal{A}(\mathbf{s}_{t})$. Depending on the action taken, the agent finds itself in a new state $\mathbf{s}_{t+1}\in\mathcal{S}$, observes the corresponding observation $\mathbf{x}_{t+1}$ and a reward $R_{t+1}\in\operatorname{\mathbb{R}}$. In the online control setting, the agent’s goal is to maximize the discounted sum of the received rewards. It may also make predictions about its environment, such as future observations’ outcomes.
@@
Williams and Zipser (1989) introduced the Real-time Recurrent Learning algorithm (RTRL) as a learning algorithm for continual recurrent learning. Instead of unrolling the recurrent dynamics back in time, RTRL computes the gradient using the most recent observation, and the gradient is calculated and carried from the last step [46]. Assuming the network parameters have not changed, this recursive form gives the exact gradient and does not suffer from the truncation bias inherent to T-BPTT. We provide a more detailed background on the BPTT and RTRL in Appendix A. In reality, the agent updates its parameters frequently, so the gradient information saved from previous time steps is stale, i.e., calculated w.r.t old parameters; yet, under the assumption of small learning rates, RTRL is known to converge [46]. These properties make RTRL ideal for online learning, but unfortunately, there is a catch: its computational complexity is quartic, of fourth order, in the size of $\mathbf{h}_{t}$, which can be prohibitively expensive. For this reason, we pursue a restricted diagonal form in this work, for which RTRL is efficient and linear in $\mathbf{h}_{t}$.
@@ 3.1 Revisiting Complex-valued Diagonal Recurrence
Since we did not impose constraints on the matrix $\mathbf{W}_{h}$, other than being diagonalizable, the eigenvalues of $\mathbf{W}_{h}$ can be complex or real numbers. Previous diagonal RNNs such as eLSTM [17], Columnar networks [18], and IndRNN [20] use only real-valued diagonal matrices. Having only real-valued diagonals assumes that the matrix $\mathbf{W}_{h}$ is symmetric. We provide a small experiment in Appendix B.2 showing that this assumption does not hold even in the simplest setting and that complex eigenvalues do arise. We also investigate whether this result can be extended beyond linear recurrence, and largely obtain a negative theortical result (see Appendix B.1 and  C ).
@@ 3.2 The RTU Parameterization
Each element of $\mathbf{h}_{t}=\mathbf{\Lambda}\mathbf{h}_{t-1}+\mathbf{W}_{x}\ \mathbf{x}_{t}\in\mathbb{R}^{2n}$ has two components $\mathbf{h}_{t}^{c_{1}},\mathbf{h}_{t}^{c_{2}}$, updated recursively:
@@
We also adopt two parameterization choices made in LRUs that showed improved performance. The first is learning logarithmic representations of the parameters rather than learning them directly: instead of learning $\mathbf{r}$ and $\boldsymbol{\theta}$, the network learns $\boldsymbol{\nu}^{\log}$ and $\boldsymbol{\theta}^{\log}$, where $\mathbf{r}\doteq\exp(-\boldsymbol{\nu})$, $\boldsymbol{\nu}=\exp(\boldsymbol{\nu}^{\log})$, and $\boldsymbol{\theta}^{\log}\doteq\log(\boldsymbol{\theta})$. This re-parametrization restricts the $\mathbf{r}$ to be $\in(0,1]$, required for stability. We found these modifications to improve stability of RTUs (see Appendix E). The second parameterization choice we adopt from LRUs is to multiply the input $(\mathbf{W}_{x}\mathbf{x}_{t})_{k}$ by a normalization factor of $\gamma_{k}={(1-r_{k}^{2})}^{1/2}$.
Putting this all together, the final formulation of RTUs is:
$\mathbf{h}^{c_{1}}_{t}$ $=\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}-\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{1}}\mathbf{x}_{t},$
$\mathbf{h}^{c_{2}}_{t}$ $=\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{2}}\mathbf{x}_{t},$ (2)
$\mathbf{h}_{t}$ $=[\mathbf{f}(\mathbf{h}_{t}^{c_{1}});\mathbf{f}(\mathbf{h}_{t}^{c_{2}})],$
@@
There are two ways to incorporate non-linearity into RTUs: inside the recurrence or after the recurrence. In the above, in Equation (2), the non-linearity is after the recurrence. These RTUs maintain the equivalence to a dense linear RNN, because the recurrence itself remains linear. We refer to this definition of RTUs as Linear RTUs, because the recurrence is linear, even though we have the ability to represent nonlinear functions by allowing for any nonlinear activation after the recurrence. We also evaluated a different variation of RTUs where the non-linearity is added to the recurrence directly. These Nonlinear RTUs are written as:
$\mathbf{h}^{c_{1}}_{t}$ $=\mathbf{f}(\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}-\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{1}}\mathbf{x}_{t}),$
$\mathbf{h}^{c_{2}}_{t}$ $=\mathbf{f}(\mathbf{g}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{2}}_{t-1}+\boldsymbol{\phi}(\boldsymbol{\nu}^{\log},\boldsymbol{\theta}^{\log})\odot\mathbf{h}^{c_{1}}_{t-1}+\boldsymbol{\gamma}\odot\mathbf{W}_{x}^{c_{2}}\mathbf{x}_{t}),$ (4)
@@ 3.3 The RTRL Update for RTUs
This section shows the RTRL updates for RTUs with more in-depth derivations in Appendix E. To keep notation simpler, we write the updates as if we are directly updating $r$ and $\theta$; the updates for $\boldsymbol{\nu}^{\log}$ and $\boldsymbol{\theta}^{\log}$ are easily obtained then using the chain rule. The full derivation is in Appendix E.2.
Consider the partial derivative with respect to $r_{1}$ for the first RTU with input $\bar{x}_{1}\doteq{(\mathbf{W}_{x}^{c_{1}}\mathbf{x}_{t})}_{1}$:
$h_{t,1}^{c_{1}}=r_{1}\cos(\theta_{1})h_{t-1,1}^{c_{1}}-r_{1}\sin(\theta_{1})h_{t-1,1}^{c_{2}}+\sqrt{(1-r_{1}^{2})}\bar{x}_{1}.$
@@
$\frac{\partial\mathcal{L}_{t}}{\partial r_{1}}=\frac{\partial\mathcal{L}_{t}}{\partial h_{t,1}^{c_{1}}}\frac{\partial h_{t,1}^{c_{1}}}{\partial r_{1}}+\frac{\partial\mathcal{L}_{t}}{\partial h_{t,1}^{c_{2}}}\frac{\partial h_{t,1}^{c_{2}}}{\partial r_{1}}.$
Since $r_{1}$ only impacts the two units in the first RTU, and derivatives w.r.t. the remaining hidden units are zero. Therefore, we just need to keep track of the vector of partial derivatives for these two components, $\mathbf{e}^{r,c_{1}}_{t}\doteq[\frac{\partial h_{t,1}^{c_{1}}}{\partial r_{1}},\ldots,\frac{\partial h_{t,n}^{c_{1}}}{\partial r_{n}}]$ and $\mathbf{e}^{r,c_{2}}_{t}\doteq[\frac{\partial h_{t,1}^{c_{2}}}{\partial r_{1}},\ldots,\frac{\partial h_{t,n}^{c_{2}}}{\partial r_{n}}]$ with recursive formulas: